The Stewart Company has $658,500 in current assets and $296,325 in current liabilities. Its initial inventory level is $164,625, and it will raise funds as additional notes payable and use them to increase inventory. How much can its short-term debt (notes payable) increase without pushing its current ratio below 2.0? Round your answer to the nearest dollar.
Answer #1

Let, Increase in short term debt and inventory be x

Current Ratio = [Current Assets + x] / [Current Liabilities + x]

2 = [$658,500 + x] / [$296,325 + x]

2 * [$296,325 + x] = [$658,500 + x]

$592,650 + 2x = $658,500 + x

2x - x = $658,500 - $592,650

x = $65,850

So, the maximum increase in short-term debt to finance inventory without lowering current ratio below 2.0 would be $65,850